What is the value of a $110 cash flow that is expected in two years, if the opportunity cost is 4 percent compounded continuously?

How do you solve this, step by step
Answer #1

We use the formula:
A=P(e)^rn
where
A=future value
P=present value
r=rate of interest
n=time period.

e=2.71828

110=P*(2.71828)^(0.04*2)

P=110/(2.71828)^(0.04*2)

=110*0.923116396

=$101.54(Approx).